Astronomy & Astrophysics manuscript no. example \begin{matrix}c \\ ◯ \end{matrix} ESO 2012 November 19, 2012

Letter to the Editor

GTC OSIRIS z-band imaging of Y dwarfs

N. N.^{1,2}, V. J. S.^{1,2}, and R. R.^{1,2,3}

1 Instituto1

e-mail: a@a 2 DD 3 EE

November 19, 2012; November 19, 2012

ABSTRACT

Aims. The aim of the project is to contribute to the characterisation of the spectral energy distribution of the coolest brown dwarfs discovered to date, the Y dwarfs. Methods. We obtained z-band far-red imaging for six Y dwarfs and i-band for one of them with the OSIRIS (Optical System for Imaging and low Resolution Integrated Spectroscopy) instrument on the 10.4-m Gran Telescopio de Canarias (GTC). Results. We detect four of six known Y dwarfs in the z-band, infer their optical-to-infrared colours, and measure their proper motions. We find a larger dispersion in the z − H colours of Y0 dwarfs than in T dwarfs.

Key words. Stars: low-mass — Stars: brown dwarfs — techniques: photometric

1. Introduction

The classical stellar sequence (?) has now been extended into very cool objects etc. etc.

2. Far-red optical imaging

2.1. Observations

The OSIRIS instrument is mounted on the 10.4-m GTC operating at the Observatory del Roque de Los Muchachos (La Palma, Canary Islands).

2.2. Data reduction and photometry

We reduced the OSIRIS Sloan z-band images in a standard manner under the IRAF environment^{1} (??).

2.3. Astrometry

We fine-tuned the astrometric calibration of the final combined science frames using IRAF and ds9 (?).

3. Discussion and conclusions

We report in this letter the first far-red optical detection of four Y0 dwarfs, the coolest brown dwarfs known to date. All four Y0 dwarfs with z-band detection have magnitudes in the range 22.8– 24.1 mag, implying z−J and z−H colours spanning 3.2–4.7 mag and 2.4-5.4 mag, respectively. We also present a lower limit of

1 IRAF is distributed by the National Optical Astronomy Observatories, which are operated by the Association of Universities for Research in Astronomy, Inc., under cooperative agreement with the National Science Foundation Fig. 1. z − H colours of late-T (open squares and diamonds; ??), UGPS J072227.51−054031.2 (open triangle; ??), and Y0 dwarfs (filled dots; this work) as a function of spectral type

z − H of 3.35 mag WISE1405+5534, a potentially peculiar Y dwarf, although its H magnitude is also a lower limit.

Acknowledgements.